If the complement of an angle is 50 degrees less than one-sixth of its supplement, find the measure of the angle.
Answer by RAY100(1637)

Let x= angle
.
Complement = 90-x
.
Supplement = 180-x
.
Given,,,,C= 1/6 S -50
.
(90-x) = (1/6) (180-x) -50
.
90-x = 30 -x/6 -50
.
110 = 5x/6
.
132 = x,,,,in degrees
.
.
check, C=90-132=-42,,,,s=180-132=48
.
-42 = 48/6 -50 = 8-50 =-42,,,,,ok
.
.